# Peeking with PEAK: Sequential, Nonparametric Composite Hypothesis Tests for Means of Multiple Data Streams

### 09 Feb 2024 v1

@@ 1 Introduction
Recent works utilizing $e$-processes for non-asymptotic anytime-valid inference fall into two groups. The first involves Chernoff-based approach (Howard et al., 2021a, ), which relies on sub-$\psi$ assumptions (such as sub-gaussian/sub-bernoulli) combined with peeling (Cappé et al.,, 2013) or mixture techniques (de la Peña et al.,, 2004) . The second involves the testing-as-betting framework (Waudby-Smith and Ramdas,, 2023). In the bounded observation setting, both empirical (Waudby-Smith and Ramdas,, 2023) and theoretical (Shekhar and Ramdas,, 2023) evidence indicate that the latter outperforms the former. However, these approaches suffer from computational inefficiency – computing confidence sequences and conducting composite tests require preforming a grid search over the hypothesis space of means, which grows in dimension as the number of data streams increases. Thus, while the current testing schemes proposed in Waudby-Smith and Ramdas, (2023) satisfy both desiderata (a), (b), they fall short of satisfying (c).
@@ 4.2 Testing Composite Convex Hypotheses
where $\gamma(m,a)=\prod_{i=1}^{t}(\mathbf{1}[A_{i}=a](c+(X_{i}-m_{a})(\hat{\mu}_{i-1}(a)-m_{a}))$ $+(1-\mathbf{1}[A_{i}=a]))$.
@@ 5 Empirical Results
Beta PEAK 88.83 $\pm$ 32.99 446.10 $\pm$ 147.08 569.67 $\pm$ 131.10 759.40 $\pm$ 191.34
Base 278.99 $\pm$ 135.71 2410.27 $\pm$ 953.47 4803.21 $\pm$ 1252.88 4994.02 $\pm$ 1620.66
Sub-B 162.59 $\pm$ 55.65 774.76 $\pm$ 295.01 1358.37 $\pm$ 413.30 1384.23 $\pm$ 409.52
Hedged 66.14 $\pm$ 26.26 355.11 $\pm$ 184.39 364.52 $\pm$ 205.41 371.09 $\pm$ 202.90
Bernoulli PEAK 97.04 $\pm$ 69.01 846.43 $\pm$ 488.32 1242.02 $\pm$ 542.04 1577.27 $\pm$ 618.99
Base 239.88 $\pm$ 93.19 2558.10 $\pm$ 739.30 4519.82 $\pm$ 917.84 4579.16 $\pm$ 910.98
Sub-B 165.87 $\pm$ 64.81 814.60 $\pm$ 347.71 1428.24 $\pm$ 455.08 1457.70 $\pm$ 450.24
Hedged 110.57 $\pm$ 71.32 1383.71 $\pm$ 1088.464 1443.82 $\pm$ 1316.45 1446.60 $\pm$ 1315.778
@@ 5.2 Multiple Streams of Data, Adaptive Sampling
Type Stopping Rule Stopping Time Correct
PEAK 189.43 $\pm$ 136.63 1.00
Beta Base 5823.40 $\pm$ 849.54 1.00
Sub-B 2413.74 $\pm$ 599.71 1.00
Hedged 383.80 $\pm$ 196.22 1.00
PEAK 267.33 $\pm$ 232.19 0.97
Bern Base 5886.56 $\pm$ 1129.41 1.00
Sub-B 3187.84 $\pm$ 3456.46 1.00
Hedged 1357.06 $\pm$ 768.06 1.00
Table 2: Stopping Times for BAI. We report averaged stopping times and one standard error over 500 simulated sample paths.
@@ 5.3 Discussion of Results
In the multi-armed case, our test performs comparably to the state of the art testing-as-betting method, Hedged, and outperforms the proposed stopping critera and Sub-B. For both the THR and BAI, the stopping times for PEAK are either the lowest or second lowest on average empirically. For THR (Table 1), the performance of our approach closely mirrors its performance in the single-arm case, where PEAK performs second-best for both high and low variance settings, and Hedged and Sub-B perform best for Beta (low variance) and Bernoulli (high variance) arms respectively.
For the best-arm identification problem, PEAK outperforms all other approaches in both settings. In the low-variance Beta setting, the average stopping time of PEAK is only half of the next smallest average stopping time, Hedged. In the high-variance Bernoulli setting, the average stopping time of PEAK is approximately $20\%$ of the next smallest average stopping time, which is also Hedged. Furthermore, in the Bernoulli setting, the correctness of our approach, is lowered to 0.97, which is closer to our 95% coverage target specified by $\alpha=0.05$.
@@
PEAK $1\times 10^{-10}$ 4.091 $\pm$ 0.198
$1\times 10^{-2}$ 3.712 $\pm$ 0.100
Hedged $5\times 10^{-3}$ 7.270 $\pm$ 0.140
$2\times 10^{-3}$ 18.334 $\pm$ 0.424
@@ A.10 Proof of Proposition 3
where $\gamma(m,a)=\prod_{i=1}^{t}\left(\mathbf{1}[A_{i}=a](c+(X_{i}-m_{a})(\hat{\mu}_{i-1}(a)-m_{a}))+(1-\mathbf{1}[A_{i}=a])\right)$.
Note that $q$ must be unique, as $\sum_{a\in W:\tilde{m}_{a}\geq q}\sum_{i=1}^{t}\gamma_{i}(a)\mathbf{1}[A_{t}=a]\frac{2q-X_{i}-\hat{\mu}_{i-1}(a)}{c+(\hat{\mu}_{i-1}(a)-q)(X_{i}-q)}$ is a monotonic function by Proposition 1. We verify that this is solution is indeed optimal through the KKT conditions (Chapt. 5, Boyd and Vandenberghe,, 2004). To see this, we first convert our problem to standard convex optimization notation:
@@
with $\theta=1/2$, 1000 grid points for each simulation, and $\lambda_{t}^{+}=\lambda_{t}^{-}=\lambda_{t}^{\text{PrPl}}$ as defined for Emp. Bern.
@@
$C_{t}(a)=\left\{\hat{\mu}_{t}(a)\pm\sqrt{\frac{\log(\frac{405.5Wt^{1.1}}{\alpha}\log(\frac{405.5Wt^{1}.1}{\delta}))}{2N_{t}(a)}}\right\}.$
@@
While all tests were done in R, we test the runtime of each approach in Python. The grid size for Hedged is 100 evenly spaced points on $[0,1]$ unless otherwise specified, and tested in the manner of the description in Table 2 on page 50 of Waudby-Smith and Ramdas, (2023). All runtimes were done locally on an Apple M2 Pro Chip, 16 Gb of RAM, with no parallelization.

### 16 Feb 2024 v2

@@ 1 Introduction
Recent works utilizing $e$-processes for non-asymptotic anytime-valid inference fall into two groups. The first involves Chernoff-based approach (Howard et al., 2021a, ), which relies on sub-$\psi$ assumptions (such as sub-gaussian/sub-bernoulli) combined with peeling (Cappé et al.,, 2013) or mixture techniques (de la Peña et al.,, 2004) . The second involves the testing-as-betting framework (Waudby-Smith and Ramdas,, 2023). In the bounded observation setting, both empirical (Waudby-Smith and Ramdas,, 2023) and theoretical (Shekhar and Ramdas,, 2023) evidence indicate that the latter outperforms the former. However, these approaches suffer from computational inefficiency – computing confidence sequences and conducting composite tests require performing a grid search over the hypothesis space of means, which grows in dimension as the number of data streams increases. Thus, while the current testing schemes proposed in Waudby-Smith and Ramdas, (2023) satisfy both desiderata (a), (b), they fall short of satisfying (c).
@@ 4.2 Testing Composite Convex Hypotheses
where $\gamma(m,a)=\prod_{i=1}^{t}(\mathbf{1}[A_{i}=a](1+(X_{i}-m_{a})(\hat{\mu}_{i-1}(a)-m_{a})/c)$ $+(1-\mathbf{1}[A_{i}=a]))$.
@@ 5 Empirical Results
Beta PEAK 77.49 $\pm$ 34.62 461.78 $\pm$ 185.64 554.52 $\pm$ 178.33 725.81 $\pm$ 218.49
Base 253.71 $\pm$ 65.39 2391.64 $\pm$ 422.35 4476.97 $\pm$ 479.41 4534.31 $\pm$ 482.94
Sub-B 199.98 $\pm$ 39.06 915.92 $\pm$ 219.18 1562.18 $\pm$ 230.75 1618.32 $\pm$ 229.90
Hedged 65.75 $\pm$ 28.73 316.74 $\pm$ 165.11 383.05 $\pm$ 161.48 479.16 $\pm$ 205.97
Bernoulli PEAK 106.75 $\pm$ 70.29 905.64 $\pm$ 479.17 1329.32 $\pm$ 513.84 1678.08$\pm$ 666.39
Base 253.83 $\pm$ 98.52 2479.43 $\pm$ 834.43 4731.87 $\pm$ 1107.71 4795.51 $\pm$ 1101.09
Sub-B 208.85 $\pm$ 64.17 944.26 $\pm$ 347.06 1624.80 $\pm$ 464.31 1686.81 $\pm$ 456.08
Hedged 104.77 $\pm$ 70.57 1212.75 $\pm$ 817.87 1582.60 $\pm$ 867.27 2241.18 $\pm$ 1092.49
@@ 5.2 Multiple Streams of Data, Adaptive Sampling
Type Stopping Rule Stopping Time
PEAK 708.52 $\pm$ 266.34
Beta Base 4686.1 $\pm$ 565.24
Sub-B 3175.3 $\pm$ 905.70
Hedged 500.12 $\pm$ 214.08
PEAK 1318.14 $\pm$ 489.29
Bern Base 4631.66 $\pm$ 896.51
Sub-B 4322.36 $\pm$ 3638.25
Hedged 1734.64 $\pm$ 858.66
Table 2: Stopping Times for BAI. We report averaged stopping times and one standard error over 100 simulated sample paths. For all simulations, regardless of method, the correct best arm was identified.
@@ 5.3 Discussion of Results
In the multi-armed case, our test performs comparably to the state of the art testing-as-betting method, Hedged, and outperforms the proposed stopping critera and Sub-B. For both the THR and BAI, the stopping times for PEAK are either the lowest or second lowest on average empirically. For THR (Table 1), the performance of our approach closely mirrors its performance in the single-arm case, where PEAK performs second-best for both high and low variance settings, and Hedged and Sub-B perform best for Beta (low variance) and Bernoulli (high variance) arms respectively. For the best-arm identification problem, the testing-as-betting methods (Hedged and PEAK) drastically outperform Base and Sub-B as a stopping criteria for the best-arm identification problem. PEAK and Hedged have the smallest or second smallest average stopping times, regardless of the distribution of arms. For the high variance, Bernoulli setting, PEAK outperforms Hedged, while Hedged outperforms PEAK in the low variance Beta setting.
@@
PEAK $1\times 10^{-10}$ 5.21 $\pm$ 0.23
$1\times 10^{-2}$ 3.72 $\pm$ 0.19
Hedged $5\times 10^{-3}$ 6.982 $\pm$ 0.17
$2.5\times 10^{-3}$ 14.50 $\pm$ 0.65
@@ A.10 Proof of Proposition 3
where $\gamma(m,a)=\prod_{i=1}^{t}\left(\mathbf{1}[A_{i}=a](1+(X_{i}-m_{a})(\hat{\mu}_{i-1}(a)-m_{a})/c)+(1-\mathbf{1}[A_{i}=a])\right)$.
We verify that this is solution is indeed optimal through the KKT conditions (Chapt. 5, Boyd and Vandenberghe,, 2004). To see this, we first convert our problem to standard convex optimization notation:
@@
with $\theta=1/2$, 100 grid points for each simulation, and $\lambda_{t}^{+}=\lambda_{t}^{-}=\lambda_{t}^{\text{PrPl}}$ as defined for the Emp. Bern. approach.
@@
$C_{t}^{\text{base}}(a)=\left\{\hat{\mu}_{t}(a)\pm\sqrt{\frac{\log(\frac{405.5Wt^{1.1}}{\alpha}\log(\frac{405.5Wt^{1}.1}{\delta}))}{2N_{t}(a)}}\right\}.$
@@
While all tests were done in R, we test the runtime of each approach in Python. The grid size for Hedged is 100, 200, and 400 evenly spaced points on $[0,1]$ (i.e. fidelity $0.01,0.005,0.0025$ respectively), and tested in the manner of the description in Table 2 on page 50 of Waudby-Smith and Ramdas, (2023). All runtimes were done locally on an Apple M2 Pro Chip, 16 Gb of RAM, with no parallelization.